OpenStudy (anonymous):

could someone help me please? 4 3/5+6 3/4=?

OpenStudy (anonymous):

4 + 6 = 10 3/5 + 3/4 lcm is 20 = 12/20 + 15/20 = 27/20 or 1 7/20 so your result is 10 + 1 7/20 = 11 7/20

OpenStudy (anonymous):

where is the 1 coming from? thats where im getting confused

OpenStudy (anonymous):

do you mean 27/20 = 1 7/20?

OpenStudy (anonymous):

divide 20 into 27 gives 1 with remainder 7 so this = 1 7/20